The unit normal to the surface x² + y² + z² = 3 at (1,1,1) is A)1/√16 (2i+2j+2k) B)1/√3(i+j+ k)

C) 1/√6(i+2j+3k)

D)1/√3(i+2j+2k)


1
Expert's answer
2021-08-17T12:30:01-0400

"x² + y² + z² = 3\\
f
(
x
,
y
,
z
)= x² + y² + z² - 3=0"

The gradient of "f\n\n(\n\nx\n\n,\n\ny\n\n,\n\nz\n\n)"  at point "x\n\n,\n\ny\n\n,\n\nz"  is a vector normal to the surface at this point.

The gradient is obtained as follows

"\u2207\nf\n(\nx\n,\ny\n,\nz\n)\n=\n(\nf_\nx\n,\nf_\ny\n,\nf_\nz\n)\n= 2x+2y+2z" at point (1,1,1) and the unit vector is

"=\frac{2,2,2}{\sqrt{2+2+2}}\\
=\frac{2}{\sqrt{8}}, \frac{2}{\sqrt{8}},\frac{2}{\sqrt{8}}\\
=\frac{1}{\sqrt{2}}, \frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}"